Please answer as soon as possible...
1) 3 graduating students who are applying for a part-time job have ages which sum up to 60. The sum of the ages of the first 2 students less the age of the third student gives 12. If the third student is 5 yrs. older than the first student, find their ages.

Answer by alicealc(293)

1st student's age = a
2nd student's age = b
3rd student's age = c

a + b + c = 60
a + b - c = 12
c = 5 + a

a + b + c = 60
a + b - c = 12
--------------- (-)
2c = 48
c = 48/2 = 24

c = 5 + a
24 = 5 + a
a = 24 - 5 = 19

a + b + c = 60
19 + b + 24 = 60
43 + b = 60
b = 60 - 43 = 17

so,
1st student's age = 19
2nd student's age = 17
3rd student's age = 24